OpenStudy (superhelp101):

Using complete sentences, explain how you would factor completely x9 - 64.

OpenStudy (superhelp101):

@zepdrix

OpenStudy (anonymous):

Medal?

OpenStudy (anonymous):

well x^9 and 64 These are both some number to the power of 3

OpenStudy (anonymous):

(x^3)^3 and 4^3

OpenStudy (anonymous):

So we can apply the formula for factoring the difference of these

OpenStudy (anonymous):

Since both terms are perfect cube roots, you first need to find the value of a=x^3 and b=4 which would look like this, ((x^3)^3-(4)^3)). Now, since the binomial can be factored using the difference of cubes formula, which is a^3-b^3=(a-b)(a^2+ab_b^2) because both terms are perfect cubes, we will get (x^3)(x^6+4x^3+16) which is our final answer.

OpenStudy (superhelp101):

Aww thx! :)

OpenStudy (anonymous):

No problem! Medal and fan please! :)